Calculus II, Spring 2019 (http://www.math.nthu.edu.tw/˜ wangwc/)

Homework Assignment for Week 01

1. Section 10.1: Problems 53, 59, 63, 67, 69, 81, 87.

2. Section 10.1: Study Definition of the limit of a sequence on page 552. Make sure that you can write down the definition correctly.

3. Section 10.2: Problems 61, 65, 71, 75.

4. Section 10.2: Study the definition of the sum of a series on page 563. See also page 1 of lecture note mse19s 190221.pdf.

5. Section 10.3: Problems 27, 31, 33, 37, 53, 55.